OpenStudy (anonymous):

Plans for a stadium are drawn on a coordinate grid. One wall lies on the line y=2x+3. A perpendicular wall passes through the point (4,-7). Write the equation of the line that contains the new wall

OpenStudy (jdoe0001):

the 1st wall lies on the line y = 2x+3 the new wall lies PERPENDICULAR to that one, and we want to know what equation would it be that makes that new wall's line so, we really have 2 walls, the old and the new one, the new perpendicular so we can say for each, we really have 2 equations one is y = 2x+3 the other we dunno it, all we know about that one is that is perpendicular to y = 2x+3, and passes through (4, -7) so 2 lines, perpendicular to each other keeping in mind that perpendicular lines have NEGATIVE RECIPROCAL slopes what would be the slope of y = 2x+3 ?

OpenStudy (anonymous):

2?

OpenStudy (jdoe0001):

yes

OpenStudy (jdoe0001):

so, if the slope of y = 2x+3, the slope of a line perpendicular to that will be RECIPROCAL of 2 1/2 NEGATIVE of that -1/2 so now we know, that it has a slope of -1/2 and passes through the point (4, -7) let us use the point-slope form of a linear equation, to plug those points in and get the equation :) point-slope form \(\large{ \implies y-y_1=m(x-x_1)\\ \quad \\ slope=m= -\cfrac{1}{2}\qquad \qquad \begin{array}{llll} &x_1&y_1\\(&4\quad ,\quad &-7) \end{array}\\ \quad \\ y-y_1=m(x-x_1)\implies y-(-7)=-\cfrac{1}{2}(x-(4))}\)

OpenStudy (jdoe0001):

just solve for "y" to get the equation

OpenStudy (anonymous):

So you would get y=-1/2x-5

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

Awesome thanks

OpenStudy (jdoe0001):

yw